# Semiregular abstract polyhedra with trivial facet stabilizer

### v1

@@ 8 Interpreting with voltage operations
In this section we want to interpret the constructions of the previous two sections as applying some well defined operation to a regular or chiral polytope. To do this we will use voltage operations. Voltage operations were introduced in [8], and they are a family of operations that include the Whytoffian constructions, duality, Petrie duality, and even some operations that modify the rank, such as the prism and the pyramid. They are very useful, because they can be applied not only to maniplexes and polytopes, but also to premaniplexes. In fact, if $\mathcal{O}$ is a voltage operation and $\mathcal{M}$ is a maniplex with symmetry type graph $\mathcal{X}$, the symmetry type graph of $\mathcal{O}(\mathcal{M})$ is a quotient of $\mathcal{O}(\mathcal{X})$ (see [8, Theorem 5.1]).
@@
An $(n,m)$-voltage operator is a voltage premaniplex $(\mathcal{Y},W^{n},\eta)$ where $\mathcal{Y}$ has rank $m$ and $W^{n}$ is the group with presentation $\langle r_{0},r_{1},\ldots r_{n-1}\mid r_{i}^{2}=(r_{i}r_{j})^{2}=1\text{ whenever }|i-j|>1\rangle$. Since the monodromy group of any $n$-premaniplex is a quotient of $W^{n}$, this group acts naturally on the set of flags of any $n$-premaniplex (on the left) by monodromies. Note that we are abusing notation by using $r_{i}$ both for an element of $W^{n}$ and an element of ${\rm Mon}(\mathcal{X})$ where $\mathcal{X}$ is a premaniplex, however this is not a problem because both elements act the same way on each flag of $\mathcal{X}$.
@@
where ${}^{i}y$ is the dart of color $i$ that starts at $y$.
Given a voltage $n$-premaniplex $(\mathcal{X},\Gamma,\xi)$ and a voltage operator $(\mathcal{Y},W^{n},\eta)$,[8, Theorem 5.4] gives us a way to construct a voltage assignment $\theta$ in $\mathcal{X}\rtimes_{\eta}\mathcal{Y}$ with values in $\Gamma$ such that $(\mathcal{X}\rtimes_{\eta}\mathcal{Y})^{\theta}$ is isomorphic to $\mathcal{X}^{\xi}\rtimes_{\eta}\mathcal{Y}$. As a corollary we get
@@
Let $\mathcal{M}$ be a regular maniplex with automorphism group $\langle\rho_{0},\rho_{1},\ldots,\rho_{n-1}\rangle$ and let $(\mathcal{Y},W^{n},\eta)$ be a voltage operator. Then $\mathcal{X}\rtimes_{\eta}\mathcal{Y}$ is isomorphic to $\mathcal{Y}^{\nu}$ where $\nu$ is the voltage assignment on $\mathcal{Y}$ we obtain from $\eta$ by replacing every occurence of each $r_{i}$ with the corresopnding $\rho_{i}$.
@@ 8.1 Family 1
Let $\mathcal{X}$ be the member of family 1 obtained from an $n$-gon. 6.3 tells us that we can construct a polytope whose symmetry type is $\mathcal{X}$ by using the voltage premaniplex in Figure 6, where $\langle\rho_{0},\ldots,\rho_{n-1}\rangle$ is the automorphism group of some non-degenerate, non-self-dual regular $n$-polytope $\mathcal{P}$ with simple Petrie polygons. Then, using 8.1 we get that $\mathcal{X}^{\xi}$ is isomorphic to $\mathcal{P}\rtimes_{\eta}\mathcal{X}$ where $(\mathcal{X},W^{n},\eta)$ is the voltage operation we get by replacing the voltage $\rho_{i}$ by $r_{i}$ (see Figure 8).
@@
In what follows we want to give some topological interpretation of what the voltage operation $\mathcal{O}:\mathcal{P}\mapsto\mathcal{P}\rtimes_{\eta}\mathcal{X}$ does. First, we need to introduce more concepts. Given a polyhedron $\mathcal{P}$ with simple Petrie-polygons, we define the petrial or Petrie dual of $\mathcal{P}$, as the polyhedron $\mathcal{P}^{\pi}$ with the same vertices and edges, but where the 2-faces correspond to the Petrie polygons of $\mathcal{P}$. We can extend this definition to all 3-maniplexes (and pre-maniplexes) as follows: Given a 3-premaniplex $\mathcal{X}$, its petrial $\mathcal{X}^{\pi}$ is defined as the premaniplex we get by removing the edges of color 0, and adding a new edge of color 0 between each flag $\Phi$ and its $02$-adjacent flag $r_{2}r_{0}\Phi$. If we denote by $\mathcal{X}^{*}$ the usual dual of $\mathcal{X}$, then $((\mathcal{X}^{*})^{\pi})^{*}$ is denoted by $\mathcal{X}^{\rm opp}$ and is known as the opposite of $\mathcal{X}$. It is easy to see that the operations dual, petrial and opposite are voltage operations. The corresponding voltage operators are ilustrated in Figure 9.
@@
Using [8, Theorem 6.1], we notice that we can interpret $\mathcal{O}$ as the result of concatenating a different operation $\mathcal{O}^{\prime}$ with the opposite operation. The operation $\mathcal{O}^{\prime}$ is given by the voltage operator $(\mathcal{Y},W^{n},\eta_{1})$, obtained by replacing the edges of color 2 in $\mathcal{X}$ by semi-edges (see Figure 10).
@@
If we want to realize an $n$-maniplex as a topological object, what we usually do is interpret each flag $\Phi$ as an $(n-1)$-simplex whose vertices are the baricenters of the faces containing that flag. We can label the vertices of this simplex according to the dimension of the face they are the baricenter of, so that the vertex labeled 0 corresponds to the vertex $\Phi_{0}$, the vertex labeled 1 corresponds to the baricenter of the edge $\Phi_{1}$, etc. Then we can number each facet of the simplex according to its opposite vertex. This means that two simplex sharing a facet labeled $i$ correspond to $i$-adjacent flags.
Figure 10 is telling us that to apply the operation $\mathcal{O}^{\prime}$ we have to replace each flag by an $n$-gon, and that $i$-adjacent flags correspond to $n$-gons sharing their $i$-th side. One way to do this topologically is as follows: We start by representing each flag as a simplex, as usual. Then, for each $i\in\{0,1,\ldots,n-1\}$ we draw a point $u_{i}$ at the baricenter of the intersection of the facets numbered $i$ and $i-1$ (reduced modulo $n$ if necessary). Then we construct an $n$-gon by drawing a sraight segment between $u_{i}$ and $u_{i+1}$ for each $i$. Notice that the edge joining $u_{i}$ to $u_{i+1}$ is contained in the facet with the label $i$. Therefore, the polygons corresponding to a pair of $i$-adjacent flags will share their $i$-th sides, just as requested.
For $n=3$, the intersection of the facet $i$ and the facet $i-1$ is exactly the vertex $i+1$, so the triangle we get by applying $\mathcal{O}^{\prime}$ is the same as the triangle representing the flag. In other words, in rank 3, the operation $\mathcal{O}^{\prime}$ is the baricentric subdivision. In rank 4 the polygons we get have a vertex in the midpoint of each edge of a particular Petrie polygon of the simplex. In Figure 11 we see an example in rank 4. Here the facets of our 4-polytope are cubes, and we illustrate the simplex corresponding to two 0-adjacent flags together with the polygons that they turn into when we apply the operation $\mathcal{O}^{\prime}$.
@@
Note that these operations can be applied to any pre-maniplex of the correct rank, but we can only be certain that the result has symmetry type in family 1 when the operated object is a regular polyhedron that is non-degenerate, non-self dual, and has simple Petrie polygons.
@@ 8.2 Family 2
Now let us consider the construction of 7.2. Remeber that this construction starts with a chiral polytope of rank $12k+$ 1 and gives us a polyhedron with $4k$-gonal facets. Using [8, Theorem 5.4] we can prove that the voltage premaniplex in Figure 7 can be obtained by applying a voltage operation $\mathcal{O}$ to the symmetry type graph of the chiral polytope $\mathcal{P}$. The voltage operator corresponding to $\mathcal{O}$ is ilustrated in Figure 12. It is similar to the operation in Figure 8, but the voltages of the edges of color 2 go from $r_{1}$ to $r_{k}$ and the edge of color 1 between the first and last flags in the cycle, has voltage $r_{0}$. We have removed the polygon in gray to keep the image cleaner.
@@
Again, this operation can be interpreted as the concatenation of another operation $\mathcal{O}^{\prime}$ and the opposite operation, which turns the edges of color 2 into semi-edges (see Figure 13).
@@
For family 2 we now give a topologic interpretation for the operation $\mathcal{O}^{\prime}$ when it turns a 3-polytope (a polyhedron) into a polyhedron with square faces. The corresponding voltage operator appears in Figure 14.
@@
One can check that the operation $\mathcal{O}^{\prime}$ in rank 3 turns every pair of 0-adjacent flags in $\mathcal{P}$ into a square. Such a pair determines an edge (1-face) and a facet of $\mathcal{P}$. Pairs of 0-adjacent flags that share a 1-face are turned into squares that share two opposite sides, while pairs that have 0-adjacent or 1-adjacent flags turn into squares that share 1 side.
A topological way to do this is as follows: Replace each edge $e$ of $\mathcal{P}$ with a cylinder split in half by a plane containing its axis. This divides the cylinder into two squares (making the ‘‘walls’’ of a digonal prism). We can label each square with one of the two facets of $\mathcal{P}$ containing $e$. One can also label each end of the cylinder with one of the vertices incident to $e$. Then, given a flag $\Phi$ of $\mathcal{P}$ there is a semicircle that corresponds to taking the cylinder corresponding to $\Phi_{1}$, taking the half labeled by $\Phi_{2}$, and taking the side of that square that corresponds to $\Phi_{0}$. Then for each flag $\Phi$ we glue the semicircle corresponding to $\Phi$ to the one corresponding to $r_{1}\Phi$. The resulting polyhedron is $\mathcal{O}^{\prime}(\mathcal{P})$.
We can interpret this operation also in the context of incidence geometries. The vertices of $\mathcal{O}^{\prime}(\mathcal{P})$ can be thought of as $\mathcal{P}_{0}\times\{0,1\}$ where $\mathcal{P}_{0}$ is the vertex-set of $\mathcal{P}$. The edges of $\mathcal{O}^{\prime}(\mathcal{P})$ can be thought of as $\mathcal{P}_{1}\times\{0,1\}$ where $\mathcal{P}_{1}$ is the edge-set of $\mathcal{P}$, together with the pairs of the form $(v,f)$ where $v$ is a vertex of $\mathcal{P}$ and $f$ is a facet of $\mathcal{P}$ incident to $v$. The edge $(e,\varepsilon)$ is incident to the vertex $(v,\delta)$ if and only if $\varepsilon=\delta$ and $e$ is incident to $v$ in $\mathcal{P}$, while the edge $(v,f)$ is incident to $(v,0)$ and $(v,1)$. The facets of $\mathcal{O}^{\prime}(\mathcal{P})$ correspond to pairs $(e,f)$ where $e$ is an edge of $\mathcal{P}$ and $f$ is a facet of $\mathcal{P}$ incident to $e$. The facet $(e,f)$ is incident to the edges $(e,0)$, $(e,1)$, $(v,f)$ and $(u,f)$, where $v$ and $u$ are the vertices of $\mathcal{P}$ incident to $e$ (see Figure 15).
@@
One can verify that the voltage operator corresponding to this operation is exactly the one in Figure 14.
Once again, the operations described in this section can be applied to all premaniplexes provided they have the correct rank, but we are only guaranteed to get a polytope with symmetry type in family 2 when the operated object is a chiral polytope with simple Petrie polygons.

### v2

@@ 8 Interpretation via voltage operations
In this section we want to interpret the constructions of the previous two sections as applying some well defined operation to a regular or chiral polytope. To do this we will use voltage operations. Voltage operations were introduced in [11], and they are a family of operations that include the Wythoffian constructions, duality, Petrie duality, and even some operations that modify the rank, such as the prism and the pyramid. They are very useful, because they can be applied not only to maniplexes and polytopes, but also to premaniplexes. In fact, if $\mathcal{O}$ is a voltage operation, $\mathcal{M}$ is a maniplex and $\mathcal{X}$ is its symmetry type graph, then the symmetry type graph of $\mathcal{O}(\mathcal{M})$ is a quotient of $\mathcal{O}(\mathcal{X})$ (see [11, Theorem 5.1]).
@@
An $(n,m)$-voltage operator is a voltage premaniplex $(\mathcal{Y},W^{n},\eta)$ where $\mathcal{Y}$ has rank $m$ and $W^{n}$ is the Coxeter group $\langle r_{0},r_{1},\ldots r_{n-1}\mid r_{i}^{2}=(r_{i}r_{j})^{2}=1\text{ whenever }|i-j|>1\rangle$. Since the monodromy group of any $n$-premaniplex is a quotient of $W^{n}$, this group acts naturally on the set of flags of any $n$-premaniplex (on the left) by monodromies. Note that we are abusing notation by using $r_{i}$ both for an element of $W^{n}$ and an element of ${\rm Mon}(\mathcal{X})$ where $\mathcal{X}$ is a premaniplex, however this is not a problem because both elements act the same way on each flag of $\mathcal{X}$.
@@
where $^{i}y$ is the dart of color $i$ that starts at $y$. In general, $\mathcal{X}\rtimes_{\eta}\mathcal{Y}$ might not be connected. However, if $\mathcal{Y}$ has a spanning tree all of whose darts have trivial voltage, and the voltages of the remaining darts generate $W^{n}$, then $\mathcal{X}\rtimes_{\eta}\mathcal{Y}$ is connected (see [11, Corollary3.10]).
Given a voltage $n$-premaniplex $(\mathcal{X},\Gamma,\xi)$ and a voltage operator $(\mathcal{Y},W^{n},\eta)$, [11, Theorem 5.4] gives us a way to construct a voltage assignment $\theta$ in $\mathcal{X}\rtimes_{\eta}\mathcal{Y}$ with values in $\Gamma$ such that $(\mathcal{X}\rtimes_{\eta}\mathcal{Y})^{\theta}$ is isomorphic to $\mathcal{X}^{\xi}\rtimes_{\eta}\mathcal{Y}$. Given a monodromy $w=r_{i_{1}}r_{i_{2}}\cdots r_{i_{\ell}}$ let us define the path $^{w}x$ as the path that starts in $x$ and follows the colors $i_{\ell},\ldots,i_{2},i_{1}$.
Theorem 8.1.
[11, Theorem 5.4] Let $(\mathcal{X},\Gamma,\xi)$ be a voltage $n$-premaniplex and $(\mathcal{Y},W^{n},\eta)$ an $(n,m)$-voltage operator. Define
$\theta(^{i}(x,y)):=\xi(^{\eta(^{i}y)}x).$
Then, $(\mathcal{X}\rtimes_{\eta}\mathcal{Y})^{\theta}$ is isomorphic to $\mathcal{X}^{\xi}\rtimes_{\eta}\mathcal{Y}$.
As a corollary we get
@@
[11, Corollary 5.6] Let $\mathcal{M}$ be a regular maniplex with automorphism group $\langle\rho_{0},\rho_{1},\ldots,\rho_{n-1}\rangle$ and let $(\mathcal{Y},W^{n},\eta)$ be a voltage operator. Then $\mathcal{M}\rtimes_{\eta}\mathcal{Y}$ is isomorphic to $\mathcal{Y}^{\nu}$ where $\nu(d)$ is obtained by writing $\eta(d)$ as a word in $\{r_{0},\ldots,r_{n-1}\}$ and replacing every occurrence of $r_{i}$ with the corresponding $\rho_{i}$.
@@ 8.1 Family 1
Let $\mathcal{X}$ be the member of family 1 obtained from an $n$-gon. ˜6.3 tells us that we can construct a polyhedron whose symmetry type is $\mathcal{X}$ by using the voltage premaniplex in Figure˜6, where $\langle\rho_{0},\ldots,\rho_{n-1}\rangle$ is the automorphism group of some non-degenerate, non-self-dual regular $n$-polytope $\mathcal{P}$ with simple Petrie polygons. Then, using ˜8.2 we get that $\mathcal{X}^{\xi}$ is isomorphic to $\mathcal{P}\rtimes_{\eta}\mathcal{X}$ where $(\mathcal{X},W^{n},\eta)$ is the voltage operation we get by replacing the voltage $\rho_{i}$ by $r_{i}$ (see Figure˜8).
@@
In what follows we want to give some topological interpretation of what the voltage operation $\mathcal{O}:\mathcal{P}\mapsto\mathcal{P}\rtimes_{\eta}\mathcal{X}$ does. First, we need to introduce more concepts. Given a polyhedron $\mathcal{P}$ with simple Petrie-polygons, we define the Petrial or Petrie dual of $\mathcal{P}$, as the polyhedron $\mathcal{P}^{\pi}$ with the same vertices and edges, but where the 2-faces correspond to the Petrie polygons of $\mathcal{P}$. We can extend this definition to all 3-maniplexes (and pre-maniplexes) as follows: Given a 3-premaniplex $\mathcal{X}$, its Petrial $\mathcal{X}^{\pi}$ is defined as the premaniplex we get by removing the edges of color 0, and adding a new edge of color 0 between each flag $\Phi$ and its $02$-adjacent flag $r_{2}r_{0}\Phi$. If we denote by $\mathcal{X}^{*}$ the usual dual of $\mathcal{X}$, then $((\mathcal{X}^{*})^{\pi})^{*}$ is denoted by $\mathcal{X}^{\rm opp}$ and is known as the opposite of $\mathcal{X}$. It is easy to see that the operations dual, Petrial and opposite are voltage operations. The corresponding voltage operators are illustrated in Figure˜9.
@@
Now we want to use the version of ˜8.1 for the concatenation of two voltage operations:
Theorem 8.3.
[11, Theorem 6.1] Let $\mathcal{X}$ be an $n$-premaniplex, $(\mathcal{Y}_{1},W^{n},\eta_{1})$ an $(n,m)$-voltage operator and $(\mathcal{Y}_{2},W^{m},\eta_{2})$ an $(m,\ell)$-voltage operator. Then
$(\mathcal{X}\rtimes_{\eta_{1}}\mathcal{Y}_{1})\rtimes_{\eta_{2}}\mathcal{Y}_{2}\cong\mathcal{X}\rtimes_{\theta}(\mathcal{Y}_{1}\rtimes_{\eta_{2}}\mathcal{Y}_{2}),$
where
$\theta(^{i}(y_{1},y_{1}))=\theta(^{\eta_{2}(^{i}y_{2})}y_{1}).$
Using ˜8.3, we notice that we can interpret $\mathcal{O}$ as the result of concatenating a different operation $\mathcal{O}^{\prime}$ with the opposite operation. The operation $\mathcal{O}^{\prime}$ is given by the voltage operator $(\mathcal{Y},W^{n},\eta_{1})$, obtained by replacing the edges of color 2 in $\mathcal{X}$ by semi-edges with the same voltage (see Figure˜10).
@@
If we want to realize an $n$-maniplex as a topological object, what we usually do is interpret each flag $\Phi$ as an (n1)(n-1)-simplex. We label the vertices of this simplex with the numbers from 0 to n1n-1 and we label each facet of the simplex according to its opposite vertex. Finally, for every two $i$-adjacent flags, we identify their facets with the label $i$ in a way that vertices with the same label are also identified. If $F$ is an $i$-face, for each flag $\Phi$ in $F$ we consider the $i$-simplex defined by the vertices of the (n1)(n-1)-simplex associated to $\Phi$ whose label is smaller or equal to $i$, and we interpret $F$ as the union of such $i$-simplices. By doing this, the vertex labeled $i$ of the simplex corresponding to the flag $\Phi$ corresponds to a fixed point in the interior of the $i$-face $(\Phi)_{i}$, which is the same for every other flag in that same $i$-face. We call this point the barycenter of $(\Phi)_{i}$.
Figure˜10 is telling us that to apply the operation $\mathcal{O}^{\prime}$ we have to replace each flag by an $n$-gon, and that $i$-adjacent flags correspond to $n$-gons sharing their $i$-th side. One way to do this topologically is as follows: We start by representing each flag as a simplex, as usual. Then, for each $i\in\{0,1,\ldots,n-1\}$ we draw a point $u_{i}$ at the barycenter of the intersection of the facets numbered $i$ and i1i-1 (reduced modulo $n$ if necessary). Then we construct an $n$-gon by drawing a straight segment between $u_{i}$ and $u_{i+1}$ for each $i$. Notice that the edge joining $u_{i}$ to $u_{i+1}$ is contained in the facet with the label $i$. Therefore, the polygons corresponding to a pair of $i$-adjacent flags will share their $i$-th sides, just as requested.
For $n=3$, the intersection of the facet $i$ and the facet i1i-1 is exactly the vertex $i+1$, so the triangle we get by applying $\mathcal{O}^{\prime}$ is the same as the triangle representing the flag. In other words, in rank 3, the operation $\mathcal{O}^{\prime}$ is the barycentric subdivision. In rank 4 the polygons we get have a vertex in the midpoint of each edge of a particular Petrie polygon of the simplex. In Figure˜11 we see an example in rank 4. Here the facets of our 4-polytope are cubes, and we illustrate the simplices corresponding to two 0-adjacent flags together with the polygons that they turn into when we apply the operation $\mathcal{O}^{\prime}$.
@@
Note that these operations can be applied to any premaniplex of the correct rank, but we can only be certain that the result has symmetry type in family 1 when the operated object is a regular polyhedron that is non-degenerate, non-self-dual, and has simple Petrie polygons.
@@ 8.2 Family 2
Now let us consider the construction of ˜7.2. Remember that this construction starts with a chiral polytope $\mathcal{P}$ of rank $2k+1$ and gives us a polyhedron $\mathcal{X}^{\xi}$ with $4k$-gonal facets. The polytope $\mathcal{P}$ can be constructed as $\mathcal{Z}^{\nu}$ where $\mathcal{Z}$ is the $(2k+1)$-premaniplex with two vertices $a$ and $b$ joined by an edge of each color from 0 to $2k$, and the voltage assignment $\nu$ is given by $\nu(^{i}a)=\tau_{i}$. Let $\mathcal{O}$ be the voltage operator in Figure˜12. It is similar to the operation in Figure˜8, but the voltages of the edges of color 2 go from $r_{1}$ to $r_{k}$ and the edge of color 1 between the first and last flags in the cycle, has voltage $r_{0}$. We have removed the polygon in gray to keep the image cleaner.
@@
Using ˜8.1 we see that the maniplex $\mathcal{X}^{\xi}$ constructed in ˜7.2 is isomorphic to the result of applying the operation $\mathcal{O}$ to the chiral polytope $\mathcal{P}\cong\mathcal{Z}^{\nu}$.
Again, the operation $\mathcal{O}$ can be interpreted as the concatenation of another operation $\mathcal{O}^{\prime}$ and the opposite operation, which turns the edges of color 2 into semi-edges (see Figure˜13).
@@
Now we would like to give a topological interpretation for the operation $\mathcal{O}^{\prime}$, however this is hard to do in high rank, so we will only do so for the case when $\mathcal{O}^{\prime}$ turns a 3-polytope (a polyhedron) into a polyhedron with square faces. The corresponding voltage operator appears in Figure˜14. Even though this particular operation does not give examples with trivial facet stabilizer (as the vertical reflection in Figure˜7 lifts), understanding this operation in rank 3 might be the first step towards understanding the ones in higher rank.
@@
One can check that the operation $\mathcal{O}^{\prime}$ in rank 3 turns every pair of 0-adjacent flags in $\mathcal{P}$ into a square. Such a pair determines an edge (1-face) and a facet of $\mathcal{P}$. Pairs of 0-adjacent flags that share a 1-face are turned into squares that share two opposite sides, while pairs that have 1-adjacent flags turn into squares that share 1 side.
A topological way to do this is as follows: Replace each edge $e$ of $\mathcal{P}$ with a cylinder split in half by a plane containing its axis. This divides the cylinder into two squares (making the ‘‘walls’’ of a digonal prism). We can label each square with one of the two facets of $\mathcal{P}$ containing $e$. One can also label each end of the cylinder with one of the vertices incident to $e$. Then, given a flag $\Phi$ of $\mathcal{P}$ there is a semicircle that corresponds to taking the cylinder corresponding to $(\Phi)_{1}$, taking the half labeled by $(\Phi)_{2}$, and taking the side of that square that corresponds to $(\Phi)_{0}$. Then for each flag $\Phi$ we glue the semicircle corresponding to $\Phi$ to the one corresponding to $r_{1}\Phi$. By doing this we get a map on the surface obtained by replacing the 1-faces of $\mathcal{P}$ by cylinders. This resulting map is the polyhedron $\mathcal{O}^{\prime}(\mathcal{P})$.
We can interpret this operation also in the context of incidence geometries. Recall that $(\mathcal{P})_{i}$ denotes the set of $i$-faces of $\mathcal{P}$. The vertices of $\mathcal{O}^{\prime}(\mathcal{P})$ can be thought of as $(\mathcal{P})_{0}\times\{0,1\}$. The edges of $\mathcal{O}^{\prime}(\mathcal{P})$ can be thought of as $(\mathcal{P})_{1}\times\{0,1\}$, together with the pairs of the form $(v,f)$ where $v$ is a vertex of $\mathcal{P}$ and $f$ is a facet of $\mathcal{P}$ incident to $v$. The edge $(e,\varepsilon)$ is incident to the vertex $(v,\delta)$ if and only if $\varepsilon=\delta$ and $e$ is incident to $v$ in $\mathcal{P}$, while the edge $(v,f)$ is incident to $(v,0)$ and $(v,1)$. The facets of $\mathcal{O}^{\prime}(\mathcal{P})$ correspond to pairs $(e,f)$ where $e$ is an edge of $\mathcal{P}$ and $f$ is a facet of $\mathcal{P}$ incident to $e$. The facet $(e,f)$ is incident to the edges $(e,0)$, $(e,1)$, $(v,f)$ and $(u,f)$, where $v$ and $u$ are the vertices of $\mathcal{P}$ incident to $e$ (see Figure˜15).
@@
One can verify that the voltage operator corresponding to this operation is exactly the one in Figure˜14.
With this interpretation it is also clear that there will always be an automorphism of $\mathcal{O}^{\prime}(\mathcal{P})$ that interchanges $(v,0)$ with $(v,1)$ for every vertex $v$, interchanges $(e,0)$ with $(e,1)$ for every edge $e$ and fixes every other element. In particular, the facet stabilizer in $\mathcal{O}^{\prime}(\mathcal{P})$ (and in $\mathcal{O}(\mathcal{P})$) is not trivial.